Given an interest rate of 9 percent, compute the year 6 future value if deposits of $3,000 and $4,000 are made in years 2 and 3, respectively, and a withdrawal of $975 is made in year 5. (Do not round intermediate calculations and round your final answer to 2 decimal places.)

Find Value in year 12:
Answer #1

Rate =9%
FV of Value in year 6 =Deposit in year 2*(1+r)^4+Deposit in year 3*(1+r)^3-Deposit Year 5*(1+r)
=3000*(1+9%)^4+4000*(1+9%)^3-975*(1+9%) =8352.1108 or 8352.11

Future Value in year 12 =8352.1108*(1+9%)^6 =14007.33